OpenStudy (anonymous):

A company's revenue (in dollars) can be modeled by the function R (p) = -15p^2 + 200p +10,000 where p is the price in dollars of the company's product. What price (in dollars) will maximize the revenue? Round to the nearest penny.

OpenStudy (apoorvk):

Differentiate and equate to zero, and find 'p' from it. That will be the value of 'p' for which the profit is maximised.

OpenStudy (anonymous):

use the theorem that when x=(-b)/2a, y gets a max/min valve. in this equation, the graph opens down, so it can have a max value where x=(-b)/2a=(-200)/(2*-15)=20/3 then substitude this x into the equation u can have a max result

OpenStudy (anonymous):

Sorry I still don't understand

OpenStudy (anonymous):

do u know that when x=(-b)/2a, y gets a max/min valve which is a theorem

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok. can u find that this equation opens down, so it have a max value

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if u know the upper, then use x=(-b)/2a to get the value of x and use that x to work out the max value